Solve. 6(2 – x) = 4x
Answer by checkley71(8403)

6(2-X)=4X
12-6X=4X
12=4X+6X
12=10X
X=12/10
X=1.2 SOLUTION
PROOF
6(2-1.2)=4*1.2
6*.8=4.8
4.8=4.8